OpenStudy (anonymous):

Explain why (–4x)0 = 1, but –4x0 = –4.

jimthompson5910 (jim_thompson5910):

anything (except 0) to the 0th power is always 1 so that's why x^0 = 1

jimthompson5910 (jim_thompson5910):

when it comes to (-4x)^0, you can think of -4x as z (ie let z = -4x) to get z^0 = 1 so that's why (-4x)^0 = 1

OpenStudy (anonymous):

\[(-4x)^0=(-1)^0(4)^0(x^0)=1\cdot1\cdot1=1\] whereas \[-4x^0=-4\cdot1=-4~~~~~~~\text{(the exponent only applies to the }x)\]